OpenStudy (peterg2001):

help please will medal Click an item in the list or group of pictures at the bottom of the problem and, holding the button down, drag it into the correct position in the answer box. Release your mouse button when the item is place. If you change your mind, drag the item to the trashcan. Click the trashcan to clear all your answers. Create a proportion using the form %/100 = is/of to represent the following word problem. Seven is what percent of 30?

OpenStudy (peterg2001):

@beastieman21

OpenStudy (peterg2001):

___/___=___/___

OpenStudy (anonymous):

1. We assume, that the number 30 is 100% - because it's the output value of the task. 2. We assume, that x is the value we are looking for. 3. If 30 is 100%, so we can write it down as 30=100%. 4. We know, that x is 7% of the output value, so we can write it down as x=7%. 5. Now we have two simple equations: 1) 30=100% 2) x=7% where left sides of both of them have the same units, and both right sides have the same units, so we can do something like that: 30/x=100%/7% 6. Now we just have to solve the simple equation, and we will get the solution we are looking for. 7. Solution for what is 7% of 30 30/x=100/7 (30/x)*x=(100/7)*x - we multiply both sides of the equation by x 30=14.2857142857*x - we divide both sides of the equation by (14.2857142857) to get x 30/14.2857142857=x 2.1=x x=2.1 now we have: 7% of 30=2.1